D | Question 13 5 pts A river flows due east at 3.0 m/s. A boat crosses the 300-m-wide river by maintaining a constant velocity of 10 m/s due north-east relative to the water. If no correction is made for the current, how far does the boat move by the time it reaches the far shore? O 173 m, upstream O 346 m, downstream O 45 m, upstream O 90 m, downstream
Answer #1

Velocity of boat, vb = 10 m/s north-east

vb = 10 cos 45 i + 10 sin 45 j

vb = (7.071 i + 7.071 j) m/s

Time taken by boat to reach the far shore,

t = 300 / 7.071 = 42.42 s

Distance, d = (resultant velocity of boat in east) * time

d = (7.071 - 3) * 42.42

d = 173 m upstream